OpenStudy (dustinrathke24):

Next-door neighbors Bob and Jim use hoses from both houses to fill Bob's swimming pool. They know that it takes 24 h using both hoses. They also know that Bob's hose, used alone, takes 20% less time than Jim's hose alone. How much time is required to fill the pool by each hose alone?

OpenStudy (welshfella):

use fractions of the pool that it takes to fill in 1 hour for the 2 hoses this is 1/24 of the pool let fraction for jims be 1 /t then for Bob this will be 0.8t so we have the equation 1/t + 1/0.8t = 1/24 solve for t

OpenStudy (welshfella):

multiply through by 25t:- 24 + 30 = t now its easy

OpenStudy (welshfella):

* by 24t